THE ELECTRONIC SPECTRUM OF TANTALUM METHYLIDYNE, TaCH

M. BARNES, \underline{DAVID A. GILLETT}, G. F. METHA^{a}, K. ATHANASSENAS and A.J. MERER, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B.C. V6T 1Z1, Canada.

Tantalum methylidyne, TaCH, has been produced in a supersonic free jet expansion in helium by the reaction between laser ablated tantalum and methan e. TaCD has been made similarly using CD_{4}. Pulsed dye laser excitation spectra from the Ω=0^{+}(X̃^{1}Σ^{+}) state of TaCH have been assigned to two electronic transitions whose upper states have Ω=0^{+}, which probably is ^{3}Π, and Ω=1. The (0, 0) bands of these states lie at 15638.6 cm^{−1} and 16398.3 cm^{−1}, respectively. Wavelength resolved fluorescence emission from the Ω=0^{+} excited state has revealed a low lying Ω=1 electronic state 3620 cm^{−1} above the ground state; this appears to be a ^{3}Δ_{1} state. Two of the X̃ state vibrational fundamentals have been determined for TaCH(TaCD): ν_{2} = 641(492) cm^{−1} and ν_{3} = 955(910) cm^{−1}. The values determined for the a^{3}Δ_{1} state were ν_{2} = 606(449) cm^{−1} and ν_{3} = 925(864) cm^{−1}. High resolution spectra are being recorded for selected bands; these will allow the molecular structure and the spectroscopic parameters to be determined accurately. The magnetic and quadrupolar hyperfine structure of 181Ta is fully resolved in the high resolution study.

aPresent address: School of Chemistry, University of Sydney, Building F11, Sydney, N.S.W. 2006, Australia